OpenStudy (anonymous):

Evaluate cot(sec^-1 (2/X))??

OpenStudy (kropot72):

The drawing below shows angle theta with \[\sec \theta=\frac{2}{x}\]  by the theorem of Pythagoras the length of the third side of the triangle is\[\sqrt{4-x ^{2}}\] Now can you find the cotangent of theta?

OpenStudy (kropot72):

Hint: In a right angled triangle the cotangent of an angle is:\[Cotangent=\frac{adjacent}{opposite}\]

OpenStudy (kropot72):

@coolgirl3517 In the drawing that I posted above, what is the side adjacent to angle theta?

OpenStudy (anonymous):

Would the answer be X divided by square root of 4-X square?

OpenStudy (kropot72):

Yes, that is the correct answer :)

OpenStudy (anonymous):

Thank you!!

OpenStudy (kropot72):

You're welcome :)